jaredc4:

What is the surface area of the right cone below? Radius 8 Slanted height 14 A. 896 units2 B. 1677 units2 C. 1767 units2 D. 2247 units

AZ:

 L is the slant height r is the radius h is the height and so it forms a right triangle and you can use Pythagorean theorem \( r^2 + h^2 = L^2\) you know that the radius is 8 and that the slant height is 14, so what is the height of the cone?

jaredc4:

alright

AZ:

\(\Large\sf Surface~ area = A=πr(r+\sqrt{h^2+r^2})\)

AZ:

Actually, there's a much better formula \(\ \Large \text{SA} = \pi r l + \pi r^2\) \( l = \text{slant height}\) \( r = \text{radius}\) and so plug in r = 8 and l = 14 \(\text{SA} = \pi(8)(14) + \pi(8^2)\) SA = ??

jaredc4:

alright thanks

jaredc4:

it dint come with a diagram sadly

AZ:

Are you sure you gave the right numbers though? The answer choices you provided don't look correct